OpenStudy (anonymous):

triangle abc is congruent to triangle def. in triangle abc, side ab measures 4 side bc x+8 and side ca 8. and def de 4 ef 3x+3 and fd 8. what equation would help solve the length of bc and ef, and why?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@paki

jimthompson5910 (jim_thompson5910):



OpenStudy (anonymous):

X+8=3x+3

jimthompson5910 (jim_thompson5910):



OpenStudy (anonymous):

And the answer would be 2.5

jimthompson5910 (jim_thompson5910):

by CPCTC, we know AB = DE BC = EF CA = FD

jimthompson5910 (jim_thompson5910):

pick on the second equation BC = EF and use substitution to get BC = EF x+8 = 3x+3

OpenStudy (anonymous):

i know the answer, i just dont know why

jimthompson5910 (jim_thompson5910):

do you see how I'm getting x+8 = 3x+3 ?

OpenStudy (anonymous):

yes, i have that, but isnt that a bit to obvious? thats all they want?

jimthompson5910 (jim_thompson5910):

well they want to know why too

OpenStudy (anonymous):

i know, but why is obvious, because if the triangles are congruent and we have matched two sides already as congruent, it leaves the last 2 sides to be congruent also

jimthompson5910 (jim_thompson5910):

yeah you just explain that the corresponding parts must be congruent (the CPCTC property) and then solving that equation gives you x once you have x, you can find the length of BC and EF

OpenStudy (anonymous):

which is 10.5

jimthompson5910 (jim_thompson5910):

correct, BC = EF = 10.5

OpenStudy (anonymous):

thats it??

jimthompson5910 (jim_thompson5910):

well you show the steps involved to getting 10.5

jimthompson5910 (jim_thompson5910):

but yeah pretty much

OpenStudy (anonymous):

nice!! thanks bro, am i supposed to do the medal thing for u or something?

jimthompson5910 (jim_thompson5910):

it's up to you and you're welcome

OpenStudy (anonymous):

i want to, but idk how